OpenStudy (anonymous):

What is the equation of a line perpendicular to the line y = -10x + 1, passing through the point (5,7)?

OpenStudy (anonymous):

slope is -10 slope of the perp is 1/10

OpenStudy (anonymous):

y-7=1/10(x-5)

OpenStudy (anonymous):

10y-70=x-5 or x-10y+65=0

OpenStudy (anonymous):

The options are in a different format. My guess was [1/10x - y = -13/2\]

OpenStudy (anonymous):

answer is just 7 pi/3